OpenStudy (anonymous):

Graph y = 2x + 3 and its inverse. Please help :)

OpenStudy (anonymous):

first you flip your x and y

OpenStudy (anonymous):

so it would be x=2y+3

OpenStudy (anonymous):

then you subtract the 3 from both sides

OpenStudy (anonymous):

which will give you x-3=2y

OpenStudy (anonymous):

are u understanding so far

OpenStudy (anonymous):

@alie-v

OpenStudy (anonymous):

Yes I am.

OpenStudy (anonymous):

k

OpenStudy (anonymous):

then you divide the 2 from the y to both sides to get the y by its self...what do you get?

OpenStudy (anonymous):

x - 3/2 = y?

OpenStudy (anonymous):

Oh wait no 2x - 3 = y right?

OpenStudy (anonymous):

almost

OpenStudy (anonymous):

2x-3/2

OpenStudy (anonymous):

Oh alright I get it

OpenStudy (anonymous):

k!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!

OpenStudy (anonymous):

So what do I do next?

OpenStudy (anonymous):

it simple just take values of x as 0,0 =+1,-1,+2,-2 and so on plot it you will get the answere